# Reliability in Semantic Segmentation: Can We Use Synthetic Data?

### v1

@@ Abstract
Assessing the reliability of perception models to covariate shifts and out-of-distribution (OOD) detection is crucial for safety-critical applications such as autonomous vehicles. By nature of the task, however, the relevant data is difficult to collect and annotate. In this paper, we challenge cutting-edge generative models to automatically synthesize data for assessing reliability in semantic segmentation. By fine-tuning Stable Diffusion [27], we perform zero-shot generation of synthetic data in OOD domains or inpainted with OOD objects. Synthetic data is employed to provide an initial assessment of pretrained segmenters, thereby offering insights into their performance when confronted with real edge cases. Through extensive experiments, we demonstrate a high correlation between the performance on synthetic data and the performance on real OOD data, showing the validity approach. Furthermore, we illustrate how synthetic data can be utilized to enhance the calibration and OOD detection capabilities of segmenters.
Figure 1: An overview of our study. Starting with in-domain data, we generate synthetic samples with the required ground truths for metrics computation. The quantitative comparison involves synthetic data vs. real data with various domain shifts and OOD objects.
$\dagger$\$$\dagger$\$footnotetext: This work was done during an internship at Valeo.ai
@@ 1 Introduction
Fueled by extensive training data, modern neural networks consistently advance accuracy across various perception benchmarks. Despite the rapid adoption of deep networks in safety-critical applications, reliability [23, 31, 33] has been an overlooked factor when designing and training these models. Recent efforts are geared toward enhancing model robustness under covariate shifts in data distributions [23, 26] and improving the model’s ability to detect the unknown [10, 36, 11]. For open-world validation, in-domain data is no longer sufficient [14]; reliable and trustworthy systems demand more rigorous testing on diverse distributions, potentially exhibiting unknown objects. However, collecting and annotating data is a challenging and costly endeavor. Data collection campaigns can be quickly overwhelmed by the growing number of out-of-distribution (OOD) conditions and objects.
While real data remains the ultimate resource for validating real-world applications, recent advances in generative networks bring hopes for a potential automatic validation framework that can gradually alleviate the need for collecting and annotating real data [17, 25]. Trained on massive datasets, state-of-the-art generative networks are regarded as foundation generalists that encompass vast semantic and domain knowledge. Furthermore, text-to-image models like Stable Diffusion [27], empowered by large vision-language models, possess the seamless prompting ability that enables zero-shot generation. Via prompting, one can intentionally generate data from any target domain with minimal effort, allowing stress-testing perception models even under extremely rare conditions. For instance, although fires or floods rarely occur in reality, self-driving cars are deemed robust and resilient; testing in such conditions is therefore crucial. Moreover, in the open world, objects not encountered during training may appear anywhere; a reliable system should be capable of recognizing its limitations and classifying such objects as unknown, referred to as OOD detection. Recent OOD detection datasets attempt to benchmark perception models with random objects found on the road [24, 12, 3, 4]; however, the number of test images and the variety of OOD objects are very limited. To tackle this challenge cost-effectively, zero-shot generative inpainting emerges as a promising tool, enabling validation across a much wider range of objects.
In this work, we exploit generative networks as a general validator, targeting multiple faces of reliability. We benchmark segmentation robustness under global covariate shifts and the ability to locally detect OOD objects. Further than that, our synthetic data can be used for calibrating pretrained models to targeted domains and for improving the OOD detection performance. To validate our proposal, we compare our synthetic data against using real test data from existing benchmarks on similar tasks, i.e. ACDC [28] and SegmentMeIfYouCan [4]. Extensive experiments demonstrate our findings as follows:
  • Synthetic data can serve as a reliable indicator of the real-life performance of segmenters under covariate shifts.
  • Pretrained segmenters can leverage synthetic data to improve confidence calibration.
  • Synthetic inpainted data is valuable for benchmarking OOD detection. Our high-quality synthetic data is featured in the official BRAVO benchmark. [^1] https://valeoai.github.io/bravo/
  • Training on inpainted OOD objects helps models become more discerning when encountering real OOD objects.
@@ 2 Related Work
Covariate Shifts. Modern machine learning models, notably deep networks, fall short in preserving their robustness and estimating their prediction confidence in the presence of covariate shifts [23, 26]. The last few years have witnessed the emergence of various benchmarks [9, 31, 6, 14, 30] that address the need for assessing models’ reliability under different distributional shifts. Hendrycks et al. [9] propose a pioneering benchmark featuring data corrupted by various synthetic perturbations such as noise, blur, and brightness, etc. Taori et al. [31] emphasize the importance of realistic shifts in reliability assessment; they highlight the disparity between natural and synthetic shifts, asserting that there is minimal to no robustness transfer from synthetic to natural distribution shifts. Sign et al. [30] studiy low-shot robustness to natural distribution shifts, highlighting the robustness properties of advanced architecture and pretraining strategies. In [5], de Jorge et al. extend beyond the standard classification setup to address the reliability problem in semantic segmentation. Most findings are largely aligned with previous works on classification; one interesting finding is on the disconnection between robustness and confidence calibration, urging more attention to calibration during segmenter design and training. On the same research line, one intriguing aspect is to study the connections between in-domain and out-of-domain robustness [21, 32, 5]; different findings suggest either positive, none or even negative correlations between ID and OOD robustness, which indeed largely depends on the type of shifts. In line with [5], we focus on segmentation as the primary research task. Different from previous works, we advocate for the use of advanced generative models to generate realistic synthetic data for testing segmenters under arbitrary covariate shifts. We align with [31] regarding the realism of testing data and argue for the rapid maturation of generative models that are progressively narrowing the gap between synthetic and real shifts. Our goal is to study whether synthetic data can be a superior choice compared to ID data in correlation studies against OOD robustness – referred to as real-shift robustness in the subsequent text to avoid confusion with OOD detection. Motivated by [5], we also attempt to address the calibration problem by leveraging our synthetic data.
@@
OOD Object Detection. In addition to robustness and calibration, the ability to detect the “unknown” is an equally important aspect, completing the reliability assessment for trustworthy systems [10, 36, 11]. In semantic segmentation, several datasets are available for evaluating OOD detection on the road, namely LostAndFound [24], StreetHazards [12] (synthetic), BDD-Anomaly [12], Fishyscapes [3] (synthetic) and SegmentMeIfYouCan [4]. Encountering and capturing images of OOD objects in real-world scenarios, without deliberately placing objects on the road, is quite uncommon. Effectively, existing datasets are limited in scale, both in terms of the number of images and the variety of object classes. In this work, we propose leveraging advanced zero-shot inpainting techniques to augment an existing segmentation dataset with inpainted OOD objects; enabling the generation of highly realistic synthetic data for testing and training OOD detection.
@@
Synthetic Data for Testing. Generative models have been exploited to generate training data either for image classification [2, 8, 29], object detection [20], or semantic segmentation [15, 38, 16, 35]. Only recently, a few works have delved into the topic of generative testing data. Li et al. [17] exploit diffusion models to realistically edit images, controlling over various object attributes. This approach enables stress-testing models and understanding their sensitivity to different attributes. Using Stable Diffusion, LANCE [25] generates counterfactual images capable of challenging any given perception model. To the extent of our knowledge, no existing works have proposed to generate testing data for segmentation reliability.
@@ 3 Reliability Under Covariate Shifts
In this section, we explore whether synthetic data can be used to assess the robustness of pretrained segmenters in the presence of covariate shifts to unseen domains. We describe the data generation process and present the benchmarking results for a wide range of segmenters on our synthetic data.
@@ 3.1 Generating images in arbitrary domains
Figure 2: Generating data with covariate shifts. Training (top) and Sampling (bottom) processes for producing the synthetic data with shifts. Note that we only use images and labels from the original domain.
Figure 3: Robustness Correlation on real vs. synthetic covariate shifts across $40$ pretrained segmenters with different architectures and network sizes. Each subplot in the second row displays the scattered mIoUs (synthetic vs. real) and the linear regression line accompanied by $95\%$ confidence intervals (CI). The subplots in the top row illustrate the Pearson Correlation Coefficients, (i) between Cityscapes mIoUs vs. real-shift mIoUs (PCC_CS  ), as well as (ii) between synthetic-shift mIoUs vs. real-shift mIoUs (PCC_Syn  ). The subplots are sorted from left to right based on domain gaps to Cityscapes, represented by PCC_CS (   ). The robustness results on synthetic data exhibit a strong correlation with those on real data, particularly in the case of the most distant shifts like ‘snow’ and ‘night’.
Our goal here is to obtain pairs of images and semantic masks, with the images belonging to visual domains for which we lack data. To this end, we leverage a pretrained text-to-image Stable Diffusion (SD) model [27], repurposed as a semantic conditioned model called ControlNet [37]. The ControlNet is trained solely on images and segmentation ground truths from the Cityscapes dataset; at train time, the text prompts are captions automatically extracted using CLIP-interrogator [1]. As a result, the model is able to perform mask-to-image generation of driving scenes while retaining the ability to steer the generation through text prompting of Stable Diffusion.
To generate synthetic data, we prompt the trained ControlNet by the concatenation of target domain descriptions and CLIP-interrogator captions obtained from Cityscapes validation images. Also, segmentation masks in the Cityscapes validation set are used to condition the generative process. Thanks to zero-shot prompting, the synthetic images are aligned with the semantic condition while displaying the visual properties of arbitrary target domains.  Figure 2 illustrates the training and generation steps. Detailed technical descriptions and more visualizations are in Appendix A.
@@ 3.2 Robustness Assessment with Synthetic Data
With the pipeline outlined in Section 3.1, one can generate synthetic data to assess the robustness of pretrained segmenters in any target unseen domains through zero-shot prompting. To quantify robustness, we employ the traditional mean Intersection-over-Union (mIoU) score, measuring the correct overlap between semantic predictions and ground-truth masks. Given that our synthetic dataset comprises pairs of segmentation masks and synthetic images, one can straightforwardly derive synthetic scores for any pretrained segmenters. We here investigate whether synthetic performance can correlatively reflect the performance on real data in target domains under covariate shifts.
@@
Experimental Setups. We address weather shifts and geographical shifts, which are often encountered in the context of autonomous driving. These covariate shifts are exhibited in the two existing real datasets: the Adverse Conditions Dataset (ACDC) [28] and the Indian Driving Dataset (IDD) [34]. We utilize those real data to quantify the quality of our synthetic data and to validate its usefulness.
@@
Synthetic data are generated by conditioning on semantic masks from the Cityscapes validation set. To prompt the ControlNet, we concatenate CLIP-interrogator’s caption with domain description following a simple template [$<$caption$>$, in $<$domain$>$] where domain is either ‘india’, ‘fog’, ‘rain’, ‘snow’ or ‘night’.
@@
For testing, we gather a collection of 40 publicly available segmenters trained only on Cityscapes, representatives of a variety of backbones, segmentation architectures, and network sizes. The full list of models is in Appendix A.
@@
Results. In Figure 3, we present our main results of this study. Our primary metric is the Pearson Correlation Coefficient (PCC) between the mIoUs on testing data vs. the mIoUs on real-shift data from ACDC’s splits or IDD. The testing data can be either the Cityscapes validation set (CS) or our synthetic data (syn); the idea is to see which testing data –whether real CS or our synthetic– correlates more with the real-shift data.
@@
We organize our results based on increasing domain gaps relative to the Cityscapes domain. The domain gaps are quantified by the Pearson correlation between Cityscapes mIoUs and real-shift mIoUs, annotated as PCC_CS and visualized as red bars   in the subplots of Figure 3. Moving from left to right, i.e. with growing domain gaps, we observe a widening discrepancy between PCC_CS and PCC_Syn. Here, PCC_Syn ( ) represents the Pearson correlation between synthetic mIoUs and real-shift mIoUs. In domains with small gaps, PCC_CS and PCC_Syn are relatively comparable. However, in domains with more adverse shifts, such as ‘snow’ and ‘night’, PCC_Syn outperforms PCC_CS significantly, exceeding PCC_CS in ‘night’ by more than double.
@@
In Figure 4, we analyze the results for the ‘night’ condition using the most robust models across different architectures, ranging from traditional convnets to recent transformer networks. We use the Semantic-FPN score as the reference to normalize the scores of other models. This normalization aims to illustrate the relative improvement in robustness concerning architecture. We rank the models from left to right based on their performance on real night data from the ACDC-night split. The consistently increasing trend of synthetic scores (   ) from left to right demonstrates a strong correlation with the ranking based on real scores. Conversely, Cityscapes scores (   ) are not indicative of night performance; that is, a higher mIoU obtained on Cityscapes does not immediately translate into a higher mIoU at night.
@@
Since synthetic data can be generated in any desired quantity, a natural question arises: how many images are sufficient? In addressing this question, we conducted experiments and presented the results in Figure 5. Our empirical finding suggests that $\sim 500$ synthetic images are adequate for a stable and reliable assessment of robustness.
Discussion. In the recent work on reliability in semantic segmentation [5], Jorge et al. systematically quantified the robustness of segmenters on real-shift data; similarly to ours, they draw comparisons from ACDC and IDD datasets. One intriguing finding in this paper is that “..the larger the domain shift, the larger the improvement brought by more recent segmentation models”, hinting at a correlation between model robustness on in-domain data and covariate-shift data. In our study, we delve deeper into this correlation, choosing to separately address different weather types instead of grouping them all together as done in [5]. In domains exhibiting small gaps to Cityscapes, such as IDD or ACDC-Fog, our conclusion aligns with [5]. However, as domain gaps increase, the discrepancy between Cityscapes mIoUs and real-shift mIoUs becomes more pronounced, resulting in poor PCC_CS scores. On the contrary, synthetic mIoUs and real-shift mIoUs exhibit a strong correlation across shifts. Our empirical study has validated that synthetic performance is a reliable indicator of model robustness in the presence of covariate shifts.
@@
Figure 5: Pearson Correlation vs. # Synthetic Samples. The use of more synthetic samples contributes to increased stability in the results. Empirical plots demonstrate that approximately $500$ samples are sufficient for a stable correlation assessment.
Figure 6: ECE improvement using synthetic data. The 40 segmenters are calibrated using either real-shift data or synthetic data. For each model, the relative ECE improvement (%) over its non-calibrated version is computed, visualized by   (synthetic shift) and   (real shift). The subplots are ranked by model robustness on real-shift data (left-to-right) and by the increased domain shift (top-to-bottom). The improvements in calibration are observed when using synthetic data, although those are much less pronounced compared to when using real data. The titles of the subplots indicate the percentage of models that showed improvement with synthetic data. In the case of weather shifts, we observe that synthetic data is particularly helpful for more robust models (i.e., the right part), while for geographical shifts, the improvement is more significant for the least robust models.

### v2

@@ Abstract
Assessing the robustness of perception models to covariate shifts and their ability to detect out-of-distribution (OOD) inputs is crucial for safety-critical applications such as autonomous vehicles. By nature of such applications, however, the relevant data is difficult to collect and annotate. In this paper, we show for the first time how synthetic data can be specifically generated to assess comprehensively the real-world reliability of semantic segmentation models. By fine-tuning Stable Diffusion [31] with only in-domain data, we perform zero-shot generation of visual scenes in OOD domains or inpainted with OOD objects. This synthetic data is employed to evaluate the robustness of pretrained segmenters, thereby offering insights into their performance when confronted with real edge cases. Through extensive experiments, we demonstrate a high correlation between the performance of models when evaluated on our synthetic OOD data and when evaluated on real OOD inputs, showing the relevance of such virtual testing. Furthermore, we demonstrate how our approach can be utilized to enhance the calibration and OOD detection capabilities of segmenters. Code and data are made public.
footnotetext: † Work done during an internship at valeo.ai
@@ 1 Introduction
Despite the rapid adoption of deep networks in safety-critical applications, reliability [26, 36, 38] has been an overlooked factor when designing and training these models. Recent efforts are geared toward enhancing model robustness under covariate shifts in data distributions [26, 30] and improving the model’s ability to detect the unknown [12, 41, 10]. For open-world validation, in-domain data is no longer sufficient [16]; reliable and trustworthy systems demand more rigorous testing on diverse distributions, potentially exhibiting unknown objects. However, data collection campaigns can be quickly overwhelmed by the growing number of out-of-distribution (OOD) objects and conditions, sometimes extreme.
Figure 1: Assessing 40 pretrained segmenters under covariate shifts. Segmentation models under scrutiny were trained on Cityscapes train set only (in-domain data). They are evaluated on (i) Cityscapes validation set, (ii) real OOD data, and (iii) proposed synthetic data. We observe a strong correlation between results on (ii) and (iii).
In this paper, we propose to leverage pretrained generative models, e.g. Stable Diffusion 1.5 (SD), to alleviate the need for real OOD data. We use SD as a general validator, targeting multiple faces of reliability:
- First, we fine-tune SD on in-domain data to enable mask conditioning; in action, zero-shot prompting generates covariate shift images for testing. Fig. 1 illustrates the generation of synthetic data in OOD domains and its use for model evaluation. By assessing 40 pretrained segmenters under covariate shifts, our first contribution demonstrates how our generated synthetic OOD benchmark can act as a powerful proxy for the real OOD benchmark ACDC [33].
- Second, we explore a similar strategy for OOD object detection assessment: we inpaint objects of unknown classes into in-domain data by injecting appropriate prompts during the zero-shot inpainting process. Inpainted images and OOD masks are used to benchmark the 40 models to see how well they can recognize OOD objects. Our results are strongly correlated to the real OOD benchmark [4]. Our high-quality synthetic data is featured in the official BRAVO benchmark.
- Third, we demonstrate the usefulness of our two synthetic OOD benchmarks for hyperparameter tuning (here model calibration) and training. We calibrate the pretrained models to targeted OOD domains using our synthetic data and validate our strategy by comparing results to using real OOD data. We also train segmenters on inpainted data for OOD detection, obtaining competitive results.
@@ 2 Related Work
Covariate Shifts. Modern machine learning models, notably deep networks, fall short in preserving their robustness and estimating their prediction confidence in the presence of covariate shifts [26, 30]. Various benchmarks [11, 36, 5, 16, 35] have emerged to address the need for assessing models’ reliability under different distributional shifts. Hendrycks et al. [11] propose a pioneering benchmark featuring data corrupted by various synthetic perturbations such as noise, blur, and brightness. Taori et al. [36] emphasize the importance of realistic shifts in reliability assessment; they highlight the disparity between natural and synthetic shifts, asserting that there is minimal to no robustness transfer from synthetic to natural distribution shifts. Sign et al. [35] study low-shot robustness to natural distribution shifts, highlighting the robustness properties of advanced architecture and pretraining strategies. In [14], de Jorge et al. extend beyond the standard classification setup to address the reliability problem in semantic segmentation, with findings largely aligned with previous works on classification; interestingly, they point to the disconnection between robustness and confidence calibration, urging more attention to calibration during segmenter design and training. On a related line, prior works study the connections between in-domain and out-of-domain robustness [23, 37, 14]; they suggest either positive, none, or even negative correlations between ID and OOD robustness, which indeed largely depends on the type of shifts. In line with [14], we focus on segmentation as the visual perception task of interest. Different from previous works, we advocate for the use of advanced generative models to generate realistic synthetic data for testing segmenters under arbitrary covariate shifts. Taori et al. [36] criticize the synthetic robustness benchmarks and advocate for using real-shift data; while we agree on the importance of realism in testing, we demonstrate that the rapid advancement of generative models now permits very meaningful virtual assessment. Our goal is to study whether synthetic data can be a superior choice compared to ID data in correlation studies against OOD robustness – referred to as real-shift robustness here to avoid confusion with OOD detection. Inspired by [14], we also calibrate models but using synthetic OOD data instead.
@@
OOD Object Detection. In addition to robustness and calibration, the ability to detect the “unknown” is equally important to assess for trustworthy systems [12, 41, 10]. In semantic segmentation, several datasets are available for evaluating OOD detection on the road, namely LostAndFound [27], StreetHazards [9] (synthetic), BDD-Anomaly [9], Fishyscapes [3] (synthetic), and SegmentMeIfYouCan [4]. Encountering and capturing images of OOD objects in real-world scenarios, without deliberately placing them on the road, is quite uncommon. Effectively, existing datasets are limited in scale, both in terms of the number of images and the variety of object classes. In this work, we propose leveraging advanced zero-shot inpainting techniques to augment an existing segmentation dataset with the insertion of OOD objects; this enables the generation of highly realistic synthetic data for testing and training OOD detection.
@@
Synthetic Data for Testing. Generative models have been exploited to create training data for image classification [2, 8, 34], object detection [22], or semantic segmentation [17, 43, 18, 40, 7]. Only recently, a few works have delved into the topic of generative testing data. Li et al. [19] exploit diffusion models to realistically edit images, controlling over various object attributes. This approach enables stress-testing models and understanding their sensitivity to different attributes. Using SD, LANCE [29] generates counterfactual images capable of challenging any given perception model. To the extent of our knowledge, no existing works have proposed to generate testing data for segmentation reliability.
@@ 3 Reliability Under Covariate Shifts
In this section, we explore whether synthetic data can be used to assess the robustness of pretrained segmenters in the presence of covariate shifts to unseen OOD domains. We describe the data generation process and present the benchmarking results for a wide range of segmenters on our synthetic data. We demonstrate the validity of the approach using domains for which a real OOD dataset exists so that we can have access to a gold standard. We stress that we do not use the OOD data at any point in the method itself. This aspect is critical for the method to be applicable to benchmark robustness in the presence of extreme or hazardous conditions.
@@ 3.1 Generating images in arbitrary domains
Figure 2: Generating data with covariate shifts. Training (left) and Sampling (right) processes for producing the synthetic data with shifts. For training, only in-domain images and masks are used. For inference, we use the in-domain masks to generate OOD images. No real OOD data is required in the framework.
Our goal here is to obtain pairs of images and semantic masks, with the images belonging to visual domains for which we lack data. To this end, we leverage a pretrained text-to-image Stable Diffusion 1.5 (SD) model [31], repurposed as a semantic-conditioned model called ControlNet [42]. ControlNet is trained solely on images and segmentation ground truths from the Cityscapes dataset; at train time, the text prompts are captions automatically extracted using CLIP-interrogator [1]. As a result, the model is able to perform mask-to-image generation of driving scenes while retaining the ability to steer the generation through text prompting of Stable Diffusion.
To generate synthetic data, we prompt a trained ControlNet by the concatenation of OOD domain descriptions and CLIP-interrogator captions obtained from Cityscapes validation images. Also, segmentation masks in Cityscapes validation set are used to condition the generative process. Thanks to zero-shot prompting, the synthetic images are aligned with the semantic condition while displaying the visual properties of arbitrary OOD domains.  Fig. 2 illustrates the training and generation steps. Detailed technical descriptions and more visualizations are in Appendix 0.A and Appendix 0.F.
@@ 3.2 Robustness Assessment with Synthetic Data
With the pipeline outlined in Sec. 3.1, one can generate synthetic data to assess the robustness of pretrained segmenters in any unseen OOD domains through zero-shot prompting. To quantify robustness, we employ the traditional mean Intersection-over-Union (mIoU) score, measuring the correct overlap between semantic predictions and ground-truth masks. Given that our synthetic dataset comprises pairs of segmentation masks and synthetic images, one can straightforwardly derive synthetic scores for any pretrained segmenters. We here investigate whether synthetic performance can faithfully reflect the performance on real data in OOD domains under covariate shifts.
Figure 3: Robustness correlation between real and synthetic covariate shifts across $40$ pretrained segmenters. The tested models, see families in bottom legend, cover different architectures and sizes. (top) Pearson Correlation Coefficients of mIoUs between Cityscapes and real-shifts (‘PCC_CS’  ), and between synthetic shifts and real ones (‘PCC_Syn’  ). (bottom) Scatter plots of synthetic vs. real mIoUs along with the linear regression line accompanied by $95\%$ confidence intervals (‘CI’). (a-e) Five types of domain shifts from Cityscapes in-domain distribution, sorted by increasing gap as assessed by decreasing PCC_CS. The robustness results on synthetic data exhibit a strong correlation with those on real data, particularly in the case of the most distant shifts like ‘snow’ and ‘night’. More details are provided in Appendix 0.D.
@@
Experimental Setups. We address weather shifts and geographical shifts, which are often encountered in autonomous driving. These covariate shifts are exhibited in two existing real datasets: the Adverse Conditions Dataset (ACDC) [33] and the Indian Driving Dataset (IDD) [39]. We utilize those real data to quantify the quality of our synthetic data and to validate its usefulness.
@@
Synthetic data are generated by conditioning on semantic masks from the Cityscapes validation set. To prompt ControlNet, we concatenate CLIP-interrogator’s caption with a domain description following a simple template [$<$caption$>$, in $<$domain$>$] where domain is either ‘india’, ‘fog’, ‘rain’, ‘snow’ or ‘night’.
@@
For testing, we gather a collection of 40 publicly available segmenters trained only on Cityscapes, representative of different backbones, segmentation architectures, and sizes. The full list of models is in Appendix 0.A.
@@
Results. In Fig. 3, we present our main results. Our primary metric is the Pearson Correlation Coefficient (PCC) between the mIoUs on testing data and on real-shift data from ACDC’s splits or from IDD. The testing data can be either the Cityscapes validation set (CS) or our synthetic data (syn); the idea is to see which testing data –whether real CS or our synthetic one– correlates more with the real-shift data. Note that in the absence of OOD data, the Cityscapes validation set acts as the closest easily available proxy and is a reasonable predictor for OOD performance as pointed out by Jorge et al. [14].
@@
We organize our results based on increasing domain gaps relative to the Cityscapes domain. The domain gaps are quantified by the Pearson correlation between Cityscapes mIoUs and real-shift mIoUs, annotated as PCC_CS and visualized as red bars   in the subplots of Fig. 3. Moving from left to right, i.e. with growing domain gaps, we observe a widening discrepancy between PCC_CS and PCC_Syn. Here, PCC_Syn ( ) represents the Pearson correlation between synthetic mIoUs and real-shift mIoUs. In domains with small gaps, PCC_CS and PCC_Syn are relatively comparable. However, in domains with more adverse shifts, such as ‘snow’ and ‘night’, PCC_Syn outperforms PCC_CS significantly, exceeding PCC_CS in ‘night’ by more than double.
@@
In Fig. 4, we analyze the results for the ‘night’ condition using the most robust models across different architectures, ranging from ConvNets to recent transformer networks. We use the Semantic-FPN score as the reference to normalize the scores of other models. This normalization aims to illustrate the relative improvement in robustness in terms of architecture. We rank the models from left to right based on their performance on real night data from the ACDC-night split. The consistently increasing trend of synthetic scores (   ) from left to right demonstrates a strong correlation with the ranking based on real scores. In contrast, Cityscapes scores (   ) are not indicative of night performance: a higher mIoU obtained on Cityscapes does not immediately translate into a higher mIoU at night.
@@
Since synthetic data can be generated in any desired quantity, a natural question arises: how many images are sufficient? In addressing this question, we conducted experiments and presented the results in Fig. 5. Our empirical finding suggests that $\sim 500$ synthetic images are adequate for a stable and reliable assessment of robustness.
@@
Figure 5: Pearson Correlation vs. # Synthetic Samples. Using more synthetic samples contributes to increased stability in the results. Empirical plots demonstrate that approximately $500$ samples are sufficient for a stable correlation assessment.
Discussion. In their recent work on reliability in semantic segmentation, Jorge et al. [14] systematically quantified the robustness of segmenters on real-shift data; similarly to ours, they draw comparisons from ACDC and IDD datasets. One intriguing finding in this paper is that “[…] the larger the domain shift, the larger the improvement brought by more recent segmentation models”, hinting at a correlation between model robustness on in-domain data and covariate-shift data; that corresponds to the CS baseline we consider here. In our study, we delve deeper into this correlation, choosing to separately address different weather types instead of grouping them all together as done in [14]. In domains exhibiting small gaps to Cityscapes, such as IDD or ACDC-Fog, our conclusion aligns with [14]. However, as domain gaps increase, the discrepancy between Cityscapes mIoUs and real-shift mIoUs becomes more pronounced, resulting in poor PCC_CS scores. On the contrary, synthetic mIoUs and real-shift mIoUs exhibit a strong correlation across shifts. Our empirical study has validated that synthetic performance is a reliable indicator of model robustness in the presence of covariate shifts.
With synthetic data, we are able to confirm observations from the literature on real data, such as (i) the robustness of transformer and ConvNext backbones and that (ii) within an architecture family, robustness correlates with the number of parameters and the robustness of the backbone.
Different generative models? The important ingredient in our pipeline is the generative model. In Tab. 1, we ablate by replacing the default SD1.5 model with other available ones: image-2-image GAN called TSIT [13], physics-based fog simulator [32], and a bigger SD variant called SDXL [28]. On ‘night’, ‘rain’, and ‘snow’, SD variants perform better than TSIT, while of note, TSIT was trained on real OOD data. The dedicated fog simulator performs very well, on par with our results using the larger model SDXL. Physics-based simulators are definitely valuable and this direction should be investigated further; however, such simulators require in-depth knowledge of OOD domains and hence are very difficult to design. Both GAN-based and physics-based approaches are limited in their scalability to many more OOD domains. Comparing SD1.5 vs. SDXL, we observe comparable results on challenging OOD domains while SDXL performs much better on ‘fog’ and ‘india’; such results hint at the future potential of stronger and better generative models in further advancing virtual testing. Unfortunately, as SDXL is much more memory-demanding, we only limit our experiments with SDXL in this particular study.
In Appendix 0.C, we report the FID scores, which measure the direct distance between synthetic and real distributions; we also extend a discussion on some limitations of our framework.
OOD expertise? OOD data? Night Rain Snow Fog India
GAN-based TSIT [13] no required 0.83 0.84 0.81 - -
Physics-based Fog Sim. [32] required no - - - 0.82 -
Ours w/ SD1.5 (default) no no 0.85 0.86 0.85 0.77 0.71
Ours w/ SDXL [28] no no 0.84 0.90 0.82 0.89 0.93
Table 1: Different generative models. Our pipeline achieves better performance (PCC with real-shifts) while does not require any OOD knowledge. Of note, it’s very easy to apply to new unseen OOD domains using our pipeline, which remains challenging for GAN-based and physics-based models.